# Transfer Learning for Paediatric Sleep Apnoea Detection Using Physiology-Guided Acoustic Models

### v1

@@ Abstract
Paediatric obstructive sleep apnoea (OSA) is clinically significant yet difficult to diagnose, as children poorly tolerate sensor-based polysomnography. Acoustic monitoring provides a non-invasive alternative for home-based OSA screening, but limited paediatric data hinders the development of robust deep learning approaches. This paper proposes a transfer learning framework that adapts acoustic models pretrained on adult sleep data to paediatric OSA detection, incorporating SpO2-based desaturation patterns to enhance model training. Using a large adult sleep dataset (157 nights) and a smaller paediatric dataset (15 nights), we systematically evaluate (i) single- versus multi-task learning, (ii) encoder freezing versus full fine-turning, and (iii) the impact of delaying SpO2 labels to better align them with the acoustics and capture physiologically meaningful features. Results show that fine-tuning with SpO2 integration consistently improves paediatric OSA detection compared with baseline models without adaptation. These findings demonstrate the feasibility of transfer learning for home-based OSA screening in children and offers its potential clinical value for early diagnosis.
@@ 1 Introduction
Obstructive sleep apnoea (OSA) is a common paediatric sleep-disordered breathing condition characterised by repeated airway obstruction, leading to intermittent hypoxia, disrupted ventilation, and fragmented sleep [1]. Its prevalence is 1–5% in the general paediatric population [2], rising to 33–61% among children with obesity [3]. Untreated OSA is associated with cardiovascular, metabolic, and neurocognitive impairments, behavioural problems, growth delay, and reduced quality of life [4].
@@
Polysomnography (PSG) is the diagnostic gold standard, but requires overnight lab-based monitoring with specialised equipment and personnel, resulting in high costs and low tolerability in children. While treatments are well established [5, 6], PSG complexity limits its utility as a universal diagnostic tool, especially for habitually snoring children at risk of OSA [7]. Consequently, over 90% of OSA cases in children remain undiagnosed [8], highlighting the need for accessible, affordable, minimally invasive diagnostics [9]. Alternative approaches, such as questionnaires [10], symptom-based scoring [11], and simplified recordings [12], lack sufficient diagnostic accuracy [13]. Machine learning (ML) shows promise for automated, objective paediatric OSA diagnosis [14], and acoustic monitoring offers a low-cost, non-invasive option for home-based OSA screening [15, 16]. However, paediatric applications remain limited due to weaker respiratory sounds, developmental physiology variability, and scarcity of labelled datasets.
OSA data collection is constrained by the slow and labour-intensive process of manual scoring. Paediatric data are particularly scarce, as children have lower tolerance for sensors, which further limits model development. Transfer learning [17] provides a practical solution by adapting models pretrained on large adult datasets to the paediatric domain. While successfully applied to sleep stage prediction from adults to children [18], its use in acoustic-based OSA detection and integration of physiological information, such as oxygen desaturation, has not yet been explored.
This study proposes transfer learning for paediatric sleep apnoea detection by adapting a convolutional neural network (CNN) trained on adult acoustic data. The framework supports both a single-task head predicting OSA and a multi-task head predicting OSA and time below a night-specific oxygen saturation (SpO2) threshold. Two adaptation strategies are investigated: freezing the encoder versus full fine-tuning. Physiological priors are incorporated by modelling the temporal delay between nasal airflow cessation and oxygen desaturation. Our contributions are threefold: (i) a transfer learning framework for acoustic-based paediatric OSA detection, (ii) evaluation of single- versus multi-task learning with different fine-tuning strategies with integration of physiological priors, and (iii) demonstration that transfer learning enables scalable, non-invasive OSA screening using smartphone-based acoustics without wearable sensors.
@@ 2 Data
Paediatric sleep data were collected by Sheffield Children’s NHS Foundation Trust with ethics approval from the UK NHS Research Ethics Committees (REC reference: 24/PR/0284). 15 children (ages 1–15 years, mean $6\pm 4$; 5 females, 10 males; mean BMI $25\pm 10$) were recruited for overnight monitoring using a Home Sleep Apnoea Test (HSAT) device (SOMNOtouch™ RESP [^1] https://somnomedics.de/en/solutions/sleep_diagnostics/polygraphy-devices/somnotouch-resp/) and concurrent smartphone audio recordings in both laboratory and home settings (participant demographics in Table 1).
The HSAT measured SpO2, heart rate, airflow, respiratory effort, snoring, sleep–wake status, and body position, with an optional thorax sensor. Audio was recorded at 16 kHz/16-bit resolution in 2-minute segments via a custom mobile app (Figure 1). Apnoea events were manually scored by a registered polysomnographic technologist [^2] https://www.brpt.org/. In total, $\sim$120 hours of data were obtained across 15 nights. Audio quality was assessed by the snore-to-non-snore ratio, while HSAT signal quality was evaluated by the percentage of missing SpO2.
The adult dataset comprised 103 participants (67 males, 36 females; mean age 45 $\pm$ 13 years; mean BMI 31 $\pm$ 7), providing 157 nights and $\sim$1094 hours of recording (see [15] for more details).
@@ 3 Methods
Fig. 2: Time delay of oxygen desaturation and Mel-frequency spectrum. Red dashed rectangles mark the 15 s oxygen desaturation window, which is shifted from 0 to longer time delays to identify the delay that best discriminates between normal and abnormal sleep.
Fig. 3: Violin plots showing time spent below SpO2 baseline as a percentage within a 15 s window versus time delay of SpO2 for non-OSA, OSA and hypopnoea segments.
@@ 3.1 Feature extraction and preprocessing
Full-night audio recordings are segmented into overlapping 30 s windows $\mathcal{S}_{i}$ with a 10 s shift. Each segment is processed via short-term Fourier transform (50 ms Hann window, 20 ms hop) to obtain a power spectrogram $P(t,f)$. A 64-filter Mel-filterbank is applied to produce log-compressed Mel-spectrograms $M(t,m)$, followed by normalisation per Mel-bin to obtain $\hat{M}_{i}\in\mathbb{R}^{T\times F}$. These features serve as input to a CNN $f_{\theta}$ for binary OSA detection $y_{i}\in\{0,1\}$, with $y_{i}=1$ segments grouped into events and aligned with HSAT annotations. For physiological integration, oxygen desaturation is quantified as the percentage of time below the night-specific SpO2 baseline (3% drop) according to AASM guideline [^3] American Academy of Sleep Medicine guideline: https://aasm.org/clinical-resources/practice-standards/practice-guidelines/ within a 15 s window for each 30 s segment (see Fig. 2).
@@ 3.2 Problem formulation
where $X_{i}^{A}\in\mathbb{R}^{T\times F}$ is a Mel spectrogram segment, $y_{i}^{A}\in\{0,1\}$ is the OSA label, and $s_{i}^{A}\in[0,1]$ is the night-specific SpO2 time percentage, $N_{A}$ is number of night recordings. Let the paediatric dataset be
@@ 3.4 Transfer learning strategy
Table 1: Participant demographics and comparison of model performance (MAE and RMSE) across patients under various conditions. SNR reflects signal quality, where higher values indicate clearer snore signals. Scored AHI is based on home sleep tests scored by qualified sleep technologists. AHI Prediction: S/M: Single-/Multi-Task Learning; NF/F/UF: Non Fine-Tuning/Freezing/Un-Freezing Encoder; FD/SD: Fixed Delay (global median time delay)/Specific Delay (night-specific time delay).
ID Sex Age BMI Dur SNR Scored S- S- S- M- M- M- M-F M-UF M-F M-UF
(hrs) (dB) AHI NF F UF NF F UF -FD -FD -SD -SD
@@ 4 EXPERIMENTAL SETUP
We first established two baselines: a single-task learning (STL) model and a multi-task learning (MTL) model trained on the adult dataset without fine-tuning, directly mapping acoustic features to OSA labels. The pretrained model jointly predicts OSA probability and SpO2 desaturation; in STL, the regression branch is omitted (Fig. 1).
@@ 4.1 Training setup
Adult models were trained for 50 epochs with a batch size of 1024 and a learning rate of $1\times 10^{-3}$ using the Adam optimiser. For paediatric adaptation, 15 nights of children’s Mel spectrograms with dual labels were fine-tuned for 20 epochs (batch size 8, learning rate $1\times 10^{-5}$, early stopping 5 epochs), balancing gradual updates and gradient noise to avoid overfitting. Models were implemented in TensorFlow Keras and trained on an NVIDIA RTX 8000 GPU (48GB).
@@ 4.2 Evaluation framework
The standard clinical measure for assessing OSA is the apnoea-hypopnoea index (AHI). This is defined as the number of apnea and hypopnea events per hour of sleep:
$\text{AHI}=\frac{N_{\text{apnea}}+N_{\text{hypopnea}}}{T_{\text{sleep}}}$ (6)
where $N_{\text{apnea}}$ and $N_{\text{hypopnea}}$ denote the number of apnoea and hypopnoea events, and $T_{\text{sleep}}$ is the total sleep time in hours. Predicted AHI was derived by aggregating the model-predicted OSA events across all segments and normalising by total sleep time. In paediatric patients, OSA severity is classified based on AHI thresholds: <<5 as normal/mild, 5–10 as moderate, and >>10 as severe [20].
We used 5-fold cross-validation on the 15-night paediatric dataset, using 12 nights for fine-tuning and 3 for testing in each fold. This ensures that training and test data come from different patients, providing cross-patient validation. Model performance was assessed using mean absolute error (MAE) and root mean square error (RMSE) between the predicted AHI and the reference AHI from the scored HSAT recordings.
@@ 5 RESULTS AND DISCUSSION
Table 1 summarises AHI prediction across model variants. Fine-tuning consistently improves performance. Without adaptation, adult STL model collapses to near-zero predictions, classifying all patients as normal and failing to capture OSA severity. In contrast, fine-tuned models yield substantially higher AHI values and better discrimination. Multi-task learning generally outperforms single-task learning across all settings, regardless of encoder freezing. This highlights the benefit of jointly modelling OSA classification and SpO2 regression.
Incorporating delayed SpO2 further reduces prediction error, with night-specific delays providing the largest gains. Among models without night-specific delays, freezing the encoder performs on par with, or better than, unfreezing, suggesting that retaining pretrained acoustic representations is advantageous when physiological alignment is limited. The best results are achieved by combining MTL with full fine-tuning and subject-specific delays (M–UF–SD), which yields the lowest MAE (2.81) and RMSE (3.86).
From a clinical perspective these improvements are meaningful. Severe OSA (AHI $\geq$ 10) warrants further evaluation, and the patient may be advised to undergo additional diagnostic procedures at a hospital. For example, for patient ID 014 (scored AHI = 21), MTL with fine-tuning correctly identified the case as severe, whereas STL and non-fine-tuned models misclassified it as normal/mild. The M–UF–SD variant (multi-task learning with unfreezing encoder and night-specific delay) demonstrates particular promise in clinical relevance. It accurately identifies normal/mild patients (AHI: 0.4–4; Prediction: 0.11–1.8), reducing false positives, and provides close estimates for moderate cases (e.g., ID 013, scored AHI 5.3 vs predicted AHI 4.46). The small difference (MAE and RMSE) indicates the model’s reliability in estimating OSA severity. While some moderate cases were underestimated (e.g., IDs 005 and 007), overall trends indicate that the model reliably captures OSA severity and could support triage decisions in paediatric screening.
A key limitation of this study is the small size of the paediatric dataset (15 patients), which constrains statistical power and may limit generalisability across broader populations. Variability in age, BMI, and signal quality further amplifies this challenge. However, this limitation directly motivates our approach: by leveraging large adult datasets for pretraining and then carefully adapting to paediatric cases, we demonstrate how transfer learning and physiological alignment can mitigate the scarcity of labelled paediatric sleep data. This strategy provides a scalable path toward developing robust paediatric OSA models in data-constrained clinical settings.
@@ 6 Conclusion
In this paper, we propose a transfer learning approach to improve the paediatric diagnosis of OSA severity by exploiting knowledge from adult apnoea detection and fine-tuning it on limited paediatric data. We systematically investigated both single- and multi-task learning frameworks, incorporating oxygen desaturation as a secondary task and considering potential time-delay cues. Experimental results demonstrate that the proposed physiology-informed transfer learning framework enhances diagnostic accuracy, supporting the feasibility of non-invasive, child-friendly smartphone-based screening for clinicians and patients. While the size of available paediatric datasets remains a limiting factor, this study illustrates how transfer learning can help mitigate data scarcity and provide a scalable route for developing clinically relevant paediatric models. Future research will extend this approach by incorporating additional low-cost physiological signals, such as accelerometer or respiratory effort data (thorax and IMU motion sensors illustrated in Fig. 1), and by validating the framework in larger paediatric cohorts to strengthen clinical translation. Taken together, this work highlights transfer learning as a practical bridge between abundant adult datasets and the pressing need for accurate, accessible diagnostic tools in paediatric sleep medicine.

### v2

@@ Abstract
Paediatric obstructive sleep apnoea (OSA) is clinically significant yet difficult to diagnose, as children poorly tolerate sensor-based polysomnography. Acoustic monitoring provides a non-invasive alternative for home-based OSA screening, but limited paediatric data hinders the development of robust deep learning approaches. This paper proposes a transfer learning framework that adapts acoustic models pretrained on adult sleep data to paediatric OSA detection, incorporating SpO2-based desaturation patterns to enhance model training. Using a large adult sleep dataset (157 nights) and a smaller paediatric dataset (15 nights), we systematically evaluate (i) single- versus multi-task learning, (ii) encoder freezing versus full fine-tuning, and (iii) the impact of delaying SpO2 labels to better align them with the acoustics and capture physiologically meaningful features. Results show that fine-tuning with SpO2 integration consistently improves paediatric OSA detection compared with baseline models without adaptation . These findings demonstrate the feasibility of transfer learning for home-based OSA screening in children and offer its potential clinical value for early diagnosis.
@@ 1 Introduction
Obstructive sleep apnoea (OSA) is a common paediatric sleep-disordered breathing condition characterised by repeated uppper-airway obstruction, leading to intermittent hypoxia, disrupted ventilation, and fragmented sleep [8]. Its prevalence is 1–5% in the general paediatric population [12], rising to 33–61% among children with obesity [16]. Untreated OSA is associated with cardiovascular, metabolic, and neurocognitive impairments, behavioural problems, growth delay, and reduced quality of life [15].
Polysomnography (PSG) is the diagnostic gold standard, but requires overnight lab-based monitoring with specialised equipment and personnel, resulting in high costs and low tolerability in children. Although effective treatments exist [10, 5], the complexity and cost of PSG limit its use as a universal diagnostic tool, particularly for children with habitual snoring who are at risk of OSA [19]. Consequently, over 90% of OSA cases in children remain undiagnosed [1], and accessible, affordable, minimally invasive diagnostics is urgently needed [17].
Alternative approaches, such as questionnaires [9], symptom-based scoring [3], and simplified recordings [11], lack sufficient diagnostic accuracy [2]. Machine learning (ML) shows promise for automated, objective paediatric OSA diagnosis [14], and acoustic monitoring offers a low-cost, non-invasive option for home-based OSA screening [18, 7]. However, paediatric applications remain limited due to weaker respiratory sounds, developmental physiology variability, and scarcity of labelled datasets.
Furthermore, OSA data collection is constrained by the slow and labour-intensive process of manual scoring. Paediatric data are particularly scarce, as children have lower tolerance for sensors, which further limits model development. Transfer learning [20] provides a practical solution by adapting models pretrained on large adult datasets to the paediatric domain. While successfully applied to sleep stage prediction from adults to children [6], its use in acoustic-based OSA detection and integration of physiological information, such as oxygen desaturation, has not yet been explored.
This study proposes transfer learning for paediatric sleep apnoea detection by adapting a convolutional neural network (CNN) trained on adult acoustic data. The framework supports both a single-task head predicting OSA and a multi-task head predicting OSA and time below a night-specific oxygen saturation (SpO2) threshold. Two adaptation strategies are investigated: freezing the encoder versus full fine-tuning. Physiological priors are incorporated by modelling the temporal delay between nasal airflow cessation and oxygen desaturation. This work makes three main contributions: (i) A transfer learning framework for acoustic-based paediatric OSA detection using adult-pretrained models, (ii) A systematic comparison of single- and multi-task learning under different fine-tuning strategies, and (iii) Integration of physiological priors via SpO2 desaturation delay modelling to improve clinically meaningful severity estimation.
@@ 2 Data
Paediatric sleep data were collected by Sheffield Children’s NHS Foundation Trust with ethics approval from the UK NHS Research Ethics Committees (REC reference: 24/PR/0284). 15 children (age range 1–15 years, mean $6\pm 4$; 5 females, 10 males) were recruited for overnight monitoring. Obesity, a key OSA risk factor, is assessed via body mass index (BMI), calculated as $\text{Weight (kg)}/\text{Height (m)}^{2}$. BMI was recorded for all participants (mean BMI $25\pm 10$). Participant demographics are summarised in Table 1.
Data were acquired using a Home Sleep Apnoea Test (HSAT) device (SOMNOtouch™ RESP [^1] https://somnomedics.de/en/solutions/sleep_diagnostics/polygraphy-devices/somnotouch-resp/), alongside concurrent smartphone audio recordings in both laboratory and home settings. The HSAT system recorded SpO2, heart rate, airflow, respiratory effort, snoring, sleep–wake status, and body position, with an optional thorax sensor. Audio was recorded at 16 kHz/16-bit resolution in 2-minute segments via a custom mobile app (Figure 1). Apnoea events were manually scored by a registered polysomnographic technologist [^2] https://www.brpt.org/. In total, $\sim$120 hours of data were obtained across 15 nights. Audio quality was assessed by the snore-to-non-snore ratio, while HSAT signal quality was evaluated by the percentage of missing SpO2.
The severity of OSA is clinically assessed using the apnoea-hypopnoea index (AHI), defined as the number of apnoea and hypopnoea events per hour of sleep. In children, OSA severity is classified as normal/mild (AHI << 5), moderate (AHI = 5–10), or severe (AHI >> 10) [4].
For comparison, the adult dataset comprised 103 participants (67 males, 36 females; mean age $45\pm 13$ years; mean BMI $31\pm 7$), contributing 157 overnight recordings and approximately 1,094 hours of data. Further details of the adult cohort and data collection protocol are reported in [18].
@@ 3.1 Feature extraction and preprocessing
Full-night audio recordings were segmented into overlapping 30 s windows $\mathcal{S}_{i}$ with a 10 s shift. Each segment is processed via short-term Fourier transform (50 ms Hann window, 20 ms hop) to obtain a power spectrogram $P(t,f)$. A 64-filter Mel-filterbank is applied to produce log-compressed Mel-spectrograms $M(t,m)$, followed by normalisation per Mel-bin to obtain $\hat{M}_{i}\in\mathbb{R}^{T\times F}$. These features serve as input to a CNN $f_{\theta}$ for binary OSA detection $y_{i}\in\{0,1\}$, with $y_{i}=1$ segments grouped into events and aligned with HSAT annotations. For physiological integration, oxygen desaturation was quantified as the percentage of time below the night-specific SpO2 baseline (3% drop) according to AASM guideline [^3] American Academy of Sleep Medicine guideline: https://aasm.org/clinical-resources/practice-standards/practice-guidelines/ within a 15 s window for each 30 s segment (see Fig. 2).
Fig. 2: Time delay of oxygen desaturation and Mel-frequency spectrum. Red dashed rectangles mark the 15-s oxygen desaturation window, which is shifted from 0 to longer time delays to identify the delay that best discriminates between normal and abnormal breathing.
@@ 3.2 Problem formulation
where $X_{i}^{A}\in\mathbb{R}^{T\times F}$ is a Mel spectrogram segment, $y_{i}^{A}\in\{0,1\}$ is the OSA label, and $s_{i}^{A}\in[0,1]$ is the night-specific SpO2 time percentage, $N_{A}$ denotes the number of night recordings. Let the paediatric dataset be
@@ 3.3 Physiological information
Fig. 3: Median - 95% CI bar plots showing time spent below SpO2 baseline as a percentage within a 15 s window versus time delay of SpO2 for non-OSA, OSA and hypopnoea segments. Median and 95% CI for non-OSA events are 0.
@@ 3.4 Transfer learning strategy
Table 1: Participant demographics and comparison of model performance (MAE and RMSE) across patients under various conditions. SNR reflects signal quality, where higher values indicate cleaner snore signals. Target AHI is based on home sleep tests scored by qualified sleep technologists. AHI Prediction: S/M: Single-/Multi-Task Learning; NF/F/UF: No Fine-Tuning/Freezing/Un-Freezing Encoder; FD/SD: Fixed Delay (global median time delay)/Specific Delay (night-specific time delay). Predicted values that are closest to the target value are in bold.
Patient Demographics and Signal Attributes Target Predicted AHI
ID Sex Age BMI Dur SNR AHI S- S- S- M- M- M- M-F M-UF M-F M-UF
(hrs) (dB) NF F UF NF F UF -FD -FD -SD -SD
@@ 4 EXPERIMENTAL SETUP
We first established two baselines: a single-task learning (STL) model and a multi-task learning (MTL) model trained on the adult dataset without fine-tuning, directly mapping acoustic features to OSA labels. The pretrained model jointly predicts OSA probability and SpO2 desaturation, with the regression branch omitted in STL (Fig. 1).
@@ 4.1 Training setup
Adult models were trained for 50 epochs with a batch size of 1024 and a learning rate of $1\times 10^{-3}$ using the Adam optimiser. For paediatric adaptation, 15 nights of children’s Mel spectrograms with dual labels were fine-tuned for 20 epochs (batch size 8, learning rate $1\times 10^{-5}$, early stopping 5 epochs), balancing gradual updates and gradient noise to avoid overfitting. Models were implemented in TensorFlow Keras and trained on an RTX 8000 GPU (48GB).
@@ 4.2 Evaluation framework
Predicted AHI was derived by aggregating the model-predicted OSA events across all segments and normalising by total sleep time. We used 5-fold cross-validation on the 15-night paediatric dataset, using 12 nights for fine-tuning (9 were used for fine-tuning and 3 were used for validation) and 3 for testing in each fold. This ensures that training and test data come from different patients, providing cross-patient validation. Model performance was assessed using mean absolute error (MAE) and root mean square error (RMSE) between the predicted AHI and the reference AHI from the scored HSAT data.
@@ 5 RESULTS AND DISCUSSION
Table 1 summarises AHI prediction performance across model variants. Fine-tuning consistently improved performance compared with models trained only on adult data across the whole data set in terms of MAE and RMSE measures. Without adaptation, the adult STL model produced near-zero predictions for most patients, effectively classifying all cases as normal and failing to reflect OSA severity. In contrast, fine-tuned models generated higher and more variable AHI estimates that better matched the reference values. Across all configurations, multi-task learning (MTL) generally outperformed single-task learning (STL), indicating that jointly modelling OSA classification and SpO2 regression provides more informative representations.
Incorporating temporally delayed SpO2 labels further reduced prediction error, with night-specific delays yielding the largest improvements. Among models that did not use night-specific delays, freezing the encoder performed comparably to, or slightly better than, full fine-tuning, suggesting that preserving pretrained acoustic representations is beneficial when physiological alignment is less precise. The overall best performance was achieved by combining MTL, full fine-tuning, and subject-specific delays (M–UF–SD), which produced the lowest MAE (2.81) and RMSE (3.86).
These reductions in AHI prediction error are clinically meaningful, as an AHI $\geq$ 10 typically indicates severe paediatric OSA requiring further diagnostic evaluation. For example, patient ID 014 (reference AHI = 21) was correctly identified as severe by the fine-tuned MTL models, whereas STL and non-adapted models misclassified this case as normal or mild. The M–UF–SD configuration also showed improved severity stratification in lower-AHI cases, producing predictions within the normal/mild range (0.11–1.8) for patients with reference AHI between 0.4 and 4, thereby reducing potential false positives. Moderate cases were generally estimated with smaller error (e.g., ID 013: reference AHI 5.3 vs. predicted 4.46), although some underestimation remained (e.g., IDs 005 and 007). Overall, the results indicate that physiology-informed transfer learning improves the reliability of AHI estimation and may support triage decisions in paediatric OSA screening.
A key limitation of this study is the small paediatric sample size (15 patients), which limits statistical power and may affect generalisability. Variability in age, BMI, and recording quality further increases inter-subject variability. Nevertheless, this constraint reflects a common challenge in paediatric sleep research and directly motivates the proposed approach. By leveraging larger adult datasets for pretraining and incorporating physiological alignment during adaptation, the framework provides a practical strategy for mitigating data scarcity in paediatric OSA modelling.
@@ 6 Conclusion
This paper presented a physiology-informed transfer learning framework for estimating paediatric OSA severity from smartphone-recorded breathing sounds. Knowledge learned from adult acoustic data was adapted to paediatric recordings through fine-tuning, with oxygen desaturation incorporated as an auxiliary physiological task. We systematically compared single- and multi-task learning formulations, evaluated encoder freezing versus full fine-tuning, and examined the effect of modelling temporal delays between acoustic events and SpO2 desaturation. Results showed that multi-task fine-tuning with subject-specific delay modelling achieved the most accurate AHI estimation, demonstrating the value of integrating physiological priors into transfer learning. These findings support the feasibility of non-invasive, child-friendly, smartphone-based screening as a complementary tool for paediatric OSA assessment.
Although the limited size of the paediatric dataset remains a constraint, the proposed framework illustrates how transfer learning can enable data-efficient development of clinically relevant models. Future work will incorporate additional low-cost physiological signals, such as accelerometry and respiratory effort (the thorax and IMU motion sensors illustrated in Fig. 1), and validate the approach in larger and more diverse paediatric cohorts to further support clinical translation.